OpenStudy (jbolyefo):

I'll fan and medal Let u = <-6, 1>, v = <-5, 2>. Find -4u + 2v.

OpenStudy (jbolyefo):

@mww

OpenStudy (jbolyefo):

@welshfella

OpenStudy (mww):

Easy stuff If you multiply a vector by a scalar the scalar is just multiplied by each element in the vector So \[k \left(\begin{matrix}a \\ b\end{matrix}\right) =\left(\begin{matrix}ka \\ kb\end{matrix}\right)\]

OpenStudy (mww):

So use this formula to convert -4u and 2v into single vectors and then do what you did to add like before.